OpenStudy (pottersheep):

1 ^ -1 = .... shouldnt it be -1 and not +1?

OpenStudy (anonymous):

nope. +1 is correct

myininaya (myininaya):

\[1^{-1}=\frac{1}{1^1}=\frac{1}{1}=1\]

OpenStudy (pottersheep):

i know its correct but i unno why :<

OpenStudy (pottersheep):

THANKSSSSSSSS =D

myininaya (myininaya):

\[a^{-n}=\frac{1}{a^n}\]

myininaya (myininaya):

\[a^{-n} \neq -a^{n}\]

OpenStudy (pottersheep):

thank you :))